OpenStudy (anonymous):

Find k given that 4, k, and k^2 - 1 are consecutive terms of a geometric sequence. Please help.

OpenStudy (anonymous):

it means that 4*r=k and k*r=k²-1, solve the system of equations

OpenStudy (anonymous):

Can you set up the equations?

OpenStudy (anonymous):

the equations are those 2 4*r=k k*r=k²-1, first equation gives r=k/4, replace in 2nd equation

OpenStudy (anonymous):

What does the r stand for?

OpenStudy (anonymous):

the ratio, look http://en.wikipedia.org/wiki/Geometric_progression

OpenStudy (anonymous):

Okay, thanks so much!!